I could use some help with the five following questions regarding epidemiology.

Moving to another question will save this response uestion1 Iculate the attack rate for person who did eat pizza Ate Ice Cream Did Not Eat Ice Cream DIARRHEATOTAL DIARRHEA 39 TOTA ATE PIZZA 52 15 14 DID NOT EAT PIZZA 40 30 50187 92 45 Moving othe estion will save this response
Answer #1

Q.1 attack rate of the pizza eatemn is= no of people eaten pizza/total no of people

total no of people ate pizza=39=14=53

total people=50=42=92

hence the attack rate of pizza eaten =53/92

Q.2 answer is 12 days as the incubation period is the time duration between the expose of the disease pathogen into the body and the the time when the first symptome of the disease is shown

hence incubation period in this case is 17-5=12 days.

Q.3. Option 1 is correct, as This graph is the type of single exposure and common vehicle outbreak type as its plotted betwen the no .of cases with respect to the time of onset in the common vehicle, that is passengers and the crew members

Q.4 Option C is correct as this class has eaten both the Pizza and the icecream

as the chances of getting infective are likely tobe:

P=39/92=0.42

P+I=39/52=0.75

I=11/15=0.73

P=Pizza

I= Ice cream

Hence P+I is most likely to be infective

Q.5.during the incubation period the symptoms of the disease try to appear